# Kaidi Zhang1, Yunmiao Zhang2, Honghai Yu2, Yan Hu1, Benyou Wang122footnotemark: 2

### v4

@@ 7 Scalability
[…]
@@
To assess the potential of data-driven scaling as a performance enhancement strategy, we investigate how varying levels of agent activation affect TwinMarket’s predictive performance. We simulate markets with 10%, 20%, 40%, and 80% of agents actively trading each day, [^5] Total 1,000 agents remain socially interactive; only a subset engage in trading. The experiments are conducted from June 15 to July 15, 2023. while keeping other conditions fixed. As shown in Figure 15, higher activation levels lead to increased trading volume and stronger collective dynamics. Figure 15 further shows that TwinMarket, even with 1,000 agents, can reproduce realistic market trends over extended periods. Together, these results reveal a clear scaling law: greater activation proportion enhances both behavioral realism and market-level alignment. This highlights TwinMarket’s ability to capture macro-level market dynamics as emergent outcomes of micro-level agent interactions.
@@ 8 Conclusion and Future Work
Moving forward, we will explore the intrinsic properties of agent trust across diverse scenarios and examine broader implications, particularly in social science and role-playing applications. Additionally, we aim to gain deeper insights into LLM agent behaviors and the fundamental parallels between LLMs and human cognition. This line of research further paves the way for studying the alignment of LLMs with humans beyond value-based considerations.
@@ Appendix A Related Work
Large language models (LLMs) excel at simulating complex human behaviors, from rational decision-making and financial market analysis to replicating behavioral biases like herding and overconfidence [3, 63, 23, 22]. Beyond behavioral simulation, they enable fine-grained modeling of individual actions and large-scale emergent phenomena [1, 2]. For instance, LLM-based multi-agent systems like TrendSim have been developed to simulate specific social media dynamics such as the spread of trending topics and the impact of poisoning attacks, incorporating features like time-aware interactions and human-like agent profiles with memory and perception modules [64]. The capacity for large-scale simulation is further demonstrated by platforms such as AgentSociety, which employs LLM-driven agents to investigate complex social issues like polarization and the effects of universal basic income by modeling agent minds with emotions, needs, and cognition [65]. Similarly, SocioVerse aims to create world models for social simulation by leveraging pools of millions of real-world user profiles to align simulations in politics, news, and economics with reality [66]. General-purpose platforms like GenSim are also emerging, supporting simulations with multi-agents and incorporating error-correction mechanisms to enhance the robustness and adaptability of customized social scenarios [67]. LLMs naturally exhibit demographic diversity, such as gender, age, and education, while engaging in rich social interactions [23, 68]. By automating text analysis, generating experimental stimuli, and designing multi-agent systems, they offer powerful tools for social science research [4, 69, 70]. Crucially, the reliability and consistency of these LLM-based simulations are active areas of research, leading to the development of specialized datasets and methods to systematically evaluate and improve how LLMs embody simulated roles and maintain persona fidelity [71]. The application of such LLM-powered agents to financial market simulation, as explored in our work, promises to capture more nuanced trader behaviors and market dynamics.
Appendix B Impact Statement
Appendix C Limitation
This research is conducted solely for academic purposes. The simulation results and models presented are not intended for real-world trading or financial decision-making.
Appendix D More Experimental Analysis
D.1 Social Network Sensitive Analysis
@@
D.2 More Ablation Study
@@
D.2.1 Impact of LLM Temperature on Agent Behavior
@@
D.2.2 Sector-Specific Dynamics: Simulation Across Different Industries
To capture the heterogeneous nature of real-world financial markets, TwinMarket simulates market dynamics across 10 aggregated industry indices (detailed in Appendix E.1.1). This ablation study assesses whether our framework generates distinct and plausible dynamics for different sectors. We analyzed key statistical properties of their simulated returns, focusing on characteristics indicative of real-world financial time series, such as fat tails (kurtosis) and volatility clustering (GARCH parameters $\alpha$ and $\beta$).
@@ D.2.2 Sector-Specific Dynamics: Simulation Across Different Industries
Industry (Index) Kurtosis GARCH-$\alpha$ GARCH-$\beta$
@@
D.2.3 Impact of Social Interaction on Agent Decision-Making
@@
D.3 Reproducibility and Stability Validation
@@
Backbone Model RMSE MAE Kurtosis Lev. Effect GARCH-$\alpha$ GARCH-$\beta$
GPT-4o (3 runs) $0.023\pm 31.13\%$ $0.026\pm 36.21\%$ $4.433\pm 20.85\%$ $0.060\pm 43.01\%$ $0.047\pm 63.17\%$ $0.780\pm 27.41\%$
Gemini-1.5-Flash (3 runs) $0.024\pm 16.34\%$ $0.020\pm 23.12\%$ $4.886\pm 8.25\%$ $0.108\pm 16.52\%$ $0.032\pm 61.67\%$ $0.897\pm 13.26\%$
@@
D.4 Emergence of Opinion Leaders
@@
D.5 Qualitative Validation of Micro-Level Behaviors
@@
D.5.1 Demonstration of Investment Analysis Approaches
@@
TwinMarket agents are capable of employing recognized investment analysis techniques, as discussed in Appendix F.1. As shown in Figure 18, an agent demonstrates Fundamental Analysis by considering valuation metrics, long-term outlook, and national strategy to justify increasing its position. Similarly, Figure 19 illustrates an agent engaging in Technical Analysis, focusing on volume-price action, moving averages, and momentum to make trading decisions, while also considering stop-loss levels for risk management.
@@
Today’s Trading Summary: Regarding TTEI (Technology and Telecom Index),tech growth stocks are currently at their lowest valuation in the past three years, with a PE ratio below 30. From a fundamental perspective, this sector already holds certain investment value. Although the leading tech stocks I own, such as SMIC and Haiguang Information, have been slightly affected by market sentiment in the short term, the long-term outlook remains positive, especially given the national strategy of promoting technological self-reliance. These companies’ growth potential is still promising, so today I decided to increase my position to 35%. As for CPEI (Chemical and Pharmaceutical Index), the pharmaceutical sector has undergone some adjustments, and valuations have fallen to a relatively reasonable range. My position in this sector is currently 30%, and I plan to continue holding and monitoring it. My total portfolio is now at 65%, which is relatively conservative, with the remaining 35% in cash, ready to gradually increase positions when better buying opportunities arise in the market.
Figure 18: Example of Fundamental Analysis.
@@
Today’s Trading Summary: Seeing the recent positive news for tech stocks, I’m planning to focus more on the TTEI (Technology and Telecom Index). I already have this index in my portfolio,and the volume-price action in the last few trading days looks solid. Notably, stocks like SMIC and Haiguang Information have broken through their 5-day moving averages, and I think there’s a good chance this momentum will continue in the short term. I’m considering adding to my position in the tech sector, aiming to raise my allocation to around 30%. That said, I need to remind myself not to get too aggressive, as other sectors in my portfolio have been performing mediocrely.The Real Estate Index (REEI) has been weak lately and has hit my stop-loss level. I plan to exit and free up some capital to invest in the tech stocks.
Figure 19: Example of Technical Analysis.
D.5.2 Exhibition of Behavioral Biases
Beyond structured analytical approaches, TwinMarket agents also manifest common behavioral tendencies and socially influenced actions. The Disposition Effect is illustrated in Figure 20, with an agent intending to sell a winning position prematurely. Figure 21 showcases Lottery Preference, where an agent favors a high-risk, speculative investment despite its poor past performance. Underdiversification is evident in Figure 22, as an agent admits to concentrating on familiar assets while neglecting others.
Figure 23 provides an example of Overconfidence, where an agent displays excessive conviction in its trading strategy. Furthermore, Figure 24 captures an instance of Herd Behavior, where an agent’s trading decision is swayed by perceived widespread forum sentiment and low market volume, overriding its typical technical approach.
@@ D.5.2 Exhibition of Behavioral Biases
Market Analysis: The central bank’s interest rate cut is definitely a major positive, and market sentiment is likely to be ignited, which could lead to a short-term rebound. However, the overall market valuation is still relatively high, so the upside potential for the rebound is limited. It’s important to be ready to take profits at the right time. News and Announcements Impact: The most important news today is the central bank’s rate cut, which directly benefits the real estate and financial sectors. These two sectors should perform well in the short term. Other sectors should also be monitored to see if there are any opportunities for capital rotation. Sector Index Analysis: FSEI (Financial Services Index): I’m currently up 5.6% on this index. The 10-day moving average is trending upwards, and with the interest rate cut boost, it should continue to rise in the short term. However, the increase may not be too large, so I plan to partially reduce my position and lock in profits once it rises another 5%.
Figure 20: Examples of Disposition Effect.
@@
Market Analysis:: The market is really volatile right now, and it feels like everyone is on the sidelines. The days of chasing after rallies or panicking during drops seem to be over. Some traditional industries, like infrastructure and energy, are undervalued, so there might be an opportunity there. Sector Index Analysis: : TSEI (Tourism Index): My lottery stock! Even though it’s been a loss so far, the 5-day moving average has been slightly picking up recently. It feels like it’s about to bounce back, so I’m adding more! It’s a gamble, but who knows, maybe I’ll turn a bike into a motorcycle!
Figure 21: Examples of Lottery Preference.
@@
Market Analysis: The market is currently in a period of consolidation, with the central bank’s interest rate cut having a significant impact on short-term market movements. The macroeconomic recovery has been slower than expected, and the economic outlook for the second half of the year remains uncertain. Overall, market sentiment is quite cautious. Sector Index Analysis: FSEI (Financial Services Index) has been declining recently. From a technical perspective, there may still be room for further downside in the short term. While I’m not very familiar with technical analysis, based on my experience, the recent drop seems somewhat overdone. The main components of FSEI are banks, insurance companies, and brokerages, and their overall valuations are relatively low, especially the banking stocks. The current PE is likely at a historical low. From a long-term perspective, there is some investment value. If it continues to fall, I may consider adding more positions.CPEI (Chemical and Pharmaceutical Index) is a stock recommended by the system. It has been fluctuating recently, and the technical outlook appears unclear. I tend to focus more on the FSEI index, which I’m more familiar with, and ignore the others.
Figure 22: Examples of Underdiversification.
@@
Today’s Trading Summary: I added more positions in MEI and CPEI, diving into Heng Rui Pharmaceutical and the manufacturing sector! Also, REEL has so many policy advantages that I couldn’t resist going all in! I’m holding onto IEEI and TTEI, hoping they continue to rise. TSEI has had a good run recently, so I decided to lock in profits and sold it! Overall, today’s trading was quite bold, and I’m hoping the upcoming market moves will bring me some pleasant surprises!
Figure 23: Examples of Overconfidencce.
@@
Post: Today’s trading was quite active. Both EREI and TTEI hit my 20% profit target, so I quickly locked in some profits and sold a portion! Although I’m a technical trader, seeing many forum discussions about the adjustment of new energy vehicle subsidy policies made me a little nervous, so I reduced my position in MEI to manage risk. Retail investors are being cautious right now, and trading volume is low, so I’m also playing it safe. I hope the market volatility stays mild—I don’t think my heart can handle too much excitement!
Figure 24: Examples of Herd Behavior.
Appendix E Technical Details I: Data Statistics
E.1 Stock Data Details
E.1.1 SSE 50 and 10 Aggregated Index

### v5

@@ 7 Scalability
Figure 10: Scaling law: Performance $\propto$ Volume.
To assess the potential of data-driven scaling as a performance enhancement strategy, we investigate how varying levels of agent activation affect TwinMarket’s predictive performance. We simulate markets with 10%, 20%, 40%, and 80% of agents actively trading each day, [^5] Total 1,000 agents remain socially interactive; only a subset engage in trading. The experiments are conducted from June 15 to July 15, 2023. while keeping other conditions fixed.
@@
This figure shows the emergent price dynamics simulated with 1,000 LLM-based agents, demonstrating realistic market fluctuations and collective behavior.
As illustrated in Figure 10, increased agent activation yields not only enhanced trading volume but, more critically, a monotonic reduction in predictive error metrics (RMSE and MAE), revealing a robust scaling relationship between market participation and model fidelity. This empirical finding aligns with theoretical predictions from collective intelligence literature that expanded agent populations engender richer information aggregation environments wherein price discovery mechanisms become increasingly robust.
Figure 11 provides further empirical validation of this scaling principle, demonstrating that TwinMarket maintains both behavioral realism and computational tractability when scaled to 1,000 agents over extended temporal horizons. Collectively, these results show that sufficiently large-scale micro-level agent interactions can generate macro-level market phenomena that exhibit the statistical regularities and emergent properties characteristic of empirical financial systems.
@@ 8 Conclusion and Future Work
The implications of this work also extend to the machine learning community. Methodologically, we demonstrate that BDI-structured LLM agents can achieve higher behavioral fidelity than traditional models, offering a blueprint for more realistic multi-agent systems. Furthermore, the framework serves as a powerful engine for generating high-fidelity synthetic data, valuable for training and evaluating downstream models, such as reinforcement learning agents.
Building on these contributions, our future work will aim to enhance the framework’s scope and realism. We will address the limitations of the current study, which focuses on the unique context of the China A-share market and employs simplified mechanics such as a single daily call auction and a zero-sum environment. Future research will involve extending the simulation to diverse market structures and incorporating more complex, continuous trading mechanisms. Such enhancements will facilitate deeper investigations into the intrinsic properties of agent trust, the fundamental parallels between LLM and human cognition, and the broader implications for social science and role-playing applications, further paving the way for studying the alignment of LLMs with humans beyond value-based considerations.
Ethics Statement
All data used in this study were acquired through lawful and ethical means, in accordance with the Fair Use principle. To protect individual privacy, no personally identifiable information was involved in the research process, and all data underwent full anonymization prior to analysis. The simulation results and models presented are intended strictly for academic use and should not be used for real-world trading or financial decision-making.
Impact Statement
Acknowledgement
This work was supported by the National Natural Science Foundation of China (No. 72495151), the Major Program of the National Fund of Philosophy and Social Science of China (No. 19ZDA105), the Shenzhen Science and Technology Program (JCYJ20220818103001002), the Shenzhen Doctoral Startup Funding (RCBS20221008093330065), the Tianyuan Fund for Mathematics of the National Natural Science Foundation of China (NSFC) (12326608), the Shenzhen Key Laboratory of Cross-Modal Cognitive Computing (Grant No. ZDSYS20230626091302006), and the Shenzhen Stability Science Program 2023.
@@ Appendix A Related Work
Large language models (LLMs) excel at simulating complex human behaviors, from rational decision-making and financial market analysis to replicating behavioral biases like herding and overconfidence [3, 68, 26, 25]. Beyond behavioral simulation, they enable fine-grained modeling of individual actions and large-scale emergent phenomena [1, 2, 69]. For instance, LLM-based multi-agent systems like TrendSim have been developed to simulate specific social media dynamics such as the spread of trending topics and the impact of poisoning attacks, incorporating features like time-aware interactions and human-like agent profiles with memory and perception modules [70]. The capacity for large-scale simulation is further demonstrated by platforms such as AgentSociety, which employs LLM-driven agents to investigate complex social issues like polarization and the effects of universal basic income by modeling agent minds with emotions, needs, and cognition [71]. Similarly, SocioVerse aims to create world models for social simulation by leveraging pools of millions of real-world user profiles to align simulations in politics, news, and economics with reality [72]. General-purpose platforms like GenSim are also emerging, supporting simulations with multi-agents and incorporating error-correction mechanisms to enhance the robustness and adaptability of customized social scenarios [73]. LLMs naturally exhibit demographic diversity, such as gender, age, and education, while engaging in rich social interactions [26, 74]. By automating text analysis, generating experimental stimuli, and designing multi-agent systems, they offer powerful tools for social science research [4, 75, 76]. Moreover, LLMs demonstrate capabilities in automating iterative research and development processes, suggesting potential for self-improving simulation frameworks [77, 78]. Crucially, the reliability and consistency of these LLM-based simulations are active areas of research, leading to the development of specialized datasets and methods to systematically evaluate and improve how LLMs embody simulated roles and maintain persona fidelity [79]. The application of such LLM-powered agents to financial market simulation, as explored in our work, promises to capture more nuanced trader behaviors and market dynamics.
Appendix B More Experimental Analysis
B.1 Social Network Sensitive Analysis
@@
B.2 More Ablation Study
@@
B.2.1 Impact of LLM Temperature on Agent Behavior
@@
B.2.2 Sector-Specific Dynamics: Simulation Across Different Industries
To capture the heterogeneous nature of real-world financial markets, TwinMarket simulates market dynamics across 10 aggregated industry indices (detailed in Appendix C.1.1). This ablation study assesses whether our framework generates distinct and plausible dynamics for different sectors. We analyzed key statistical properties of their simulated returns, focusing on characteristics indicative of real-world financial time series, such as fat tails (kurtosis) and volatility clustering (GARCH parameters $\alpha$ and $\beta$).
@@ B.2.2 Sector-Specific Dynamics: Simulation Across Different Industries
Industry (Index) Kurtosis $\alpha GARCH$-$\beta$
@@
B.2.3 Impact of Social Interaction on Agent Decision-Making
@@
B.2.4 Granular Analysis of the BDI Framework
To provide a more granular analysis of the Belief-Desire-Intention (BDI) framework’s contribution, we conducted a targeted ablation study on its core cognitive components. We evaluated two additional variants against the full model:
  • TwinMarket w/o Belief Update: In this variant, agents maintain their initial belief state throughout the simulation. The dynamic belief adaptation mechanism is disabled, preventing agents from updating their views based on new market or social information.
  • TwinMarket w/o Desire: Here, agents skip the proactive information-seeking step. They do not generate queries for external information and make decisions based solely on their pre-existing knowledge and past experiences.
We conducted a month-long simulation from 2023-06-15 to 2023-07-15 with 100 agents powered by GPT-4o. The results, summarized in Table 9, show that removing either component leads to a significant degradation in predictive accuracy. The performance drop in the ‘w/o Belief Update‘ variant highlights the necessity of dynamic adaptation for capturing market realism. Similarly, the weaker performance of the ‘w/o Desire‘ variant underscores the importance of goal-driven, information-seeking behavior. These findings confirm that both the adaptive belief mechanism and the proactive desire component are indispensable for the framework’s ability to replicate complex market dynamics.
Table 9: Granular ablation study on the core components of the BDI framework.
Model Variant RMSE $\downarrow$ MAE $\downarrow$
TwinMarket (Full BDI) 0.0158 0.0143
TwinMarket w/o Belief Update 0.0342 0.0298
TwinMarket w/o Desire 0.0532 0.0443
B.3 Reproducibility and Stability Validation
@@
Backbone Model RMSE MAE Kurtosis Lev. Effect $\alpha GARCH$-$\beta$
GPT-4o (3 runs) $023\pm 31.13\% 0.0$$026\pm 36.21\%$ $4.433\pm 20.85\%$ $060\pm 43.01\% 0.0$$047\pm 63.17\%$ $0.780\pm 27.41\%$
Gemini-1.5-Flash (3 runs) $024\pm 16.34\% 0.0$$020\pm 23.12\%$ $4.886\pm 8.25\%$ $108\pm 16.52\% 0.0$$032\pm 61.67\%$ $0.897\pm 13.26\%$
@@
B.4 Emergence of Opinion Leaders
@@
B.5 Qualitative Validation of Micro-Level Behaviors
@@
B.5.1 Demonstration of Investment Analysis Approaches
@@
TwinMarket agents are capable of employing recognized investment analysis techniques, as discussed in Appendix D.1. As shown in Figure 14(a), an agent demonstrates Fundamental Analysis by considering valuation metrics, long-term outlook, and national strategy to justify increasing its position. Similarly, Figure 14(b) illustrates an agent engaging in Technical Analysis, focusing on volume-price action, moving averages, and momentum to make trading decisions, while also considering stop-loss levels for risk management.
{examplebox}
@@
(a) Fundamental Analysis Example Today’s Trading Summary: Regarding TTEI (Technology and Telecom Index), tech growth stocks are currently at their lowest valuation in the past three years, with a PE ratio below 30. From a fundamental perspective, this sector already holds certain investment value. Although the leading tech stocks I own, such as SMIC and Haiguang Information, have been slightly affected by market sentiment in the short term, the long-term outlook remains positive, especially given the national strategy of promoting technological self-reliance. These companies’ growth potential is still promising, so today I decided to increase my position to 35%.
As for CPEI (Chemical and Pharmaceutical Index), the pharmaceutical sector has undergone some adjustments, and valuations have fallen to a relatively reasonable range. My position in this sector is currently 30%, and I plan to continue holding and monitoring it. My total portfolio is now at 65%, which is relatively conservative, with the remaining 35% in cash, ready to gradually increase positions when better buying opportunities arise in the market.
{examplebox}
@@
(b) Technical Analysis Example Today’s Trading Summary: Seeing the recent positive news for tech stocks, I’m planning to focus more on the TTEI (Technology and Telecom Index). I already have this index in my portfolio, and the volume-price action in the last few trading days looks solid. Notably, stocks like SMIC and Haiguang Information have broken through their 5-day moving averages, and I think there’s a good chance this momentum will continue in the short term. I’m considering adding to my position in the tech sector, aiming to raise my allocation to around 30%.
That said, I need to remind myself not to get too aggressive, as other sectors in my portfolio have been performing mediocrely. The Real Estate Index (REEI) has been weak lately and has hit my stop-loss level. I plan to exit and free up some capital to invest in the tech stocks.
Figure 14: Examples of trading analysis methods: (a) fundamental analysis based on valuations, and (b) technical analysis based on price movements.
B.5.2 Exhibition of Behavioral Biases
Beyond structured analytical approaches, TwinMarket agents also manifest common behavioral tendencies and socially influenced actions.
The Disposition Effect is illustrated in Figure 15, with an agent intending to sell a winning position prematurely. {examplebox}Example of Disposition Effect Market Analysis: The central bank’s interest rate cut is definitely a major positive, and market sentiment is likely to be ignited, which could lead to a short-term rebound. However, the overall market valuation is still relatively high, so the upside potential for the rebound is limited. It’s important to be ready to take profits at the right time.
News and Announcements Impact: The most important news today is the central bank’s rate cut, which directly benefits the real estate and financial sectors. These two sectors should perform well in the short term. Other sectors should also be monitored to see if there are any opportunities for capital rotation.
@@ B.5.2 Exhibition of Behavioral Biases
Sector Index Analysis: FSEI (Financial Services Index): I’m currently up 5.6% on this index. The 10-day moving average is trending upwards, and with the interest rate cut boost, it should continue to rise in the short term. However, the increase may not be too large, so I plan to partially reduce my position and lock in profits once it rises another 5%.
Figure 15: Example of Disposition Effect.
Figure 16 showcases Lottery Preference, where an agent favors a high-risk, speculative investment despite its poor past performance.
{examplebox}
@@
Example of Lottery Preference Market Analysis: The market is really volatile right now, and it feels like everyone is on the sidelines. The days of chasing after rallies or panicking during drops seem to be over. Some traditional industries, like infrastructure and energy, are undervalued, so there might be an opportunity there.
Sector Index Analysis: TSEI (Tourism Index): My lottery stock! Even though it’s been a loss so far, the 5-day moving average has been slightly picking up recently. It feels like it’s about to bounce back, so I’m adding more! It’s a gamble, but who knows, maybe I’ll turn a bike into a motorcycle!
Figure 16: Example of Lottery Preference.
Underdiversification is evident in Figure 17, as an agent admits to concentrating on familiar assets while neglecting others.
{examplebox}
Example of Underdiversification Market Analysis: The market is currently in a period of consolidation, with the central bank’s interest rate cut having a significant impact on short-term market movements. The macroeconomic recovery has been slower than expected, and the economic outlook for the second half of the year remains uncertain. Overall, market sentiment is quite cautious.
@@
Sector Index Analysis: FSEI (Financial Services Index) has been declining recently. From a technical perspective, there may still be room for further downside in the short term. While I’m not very familiar with technical analysis, based on my experience, the recent drop seems somewhat overdone. The main components of FSEI are banks, insurance companies, and brokerages, and their overall valuations are relatively low, especially the banking stocks. The current PE is likely at a historical low. From a long-term perspective, there is some investment value. If it continues to fall, I may consider adding more positions. CPEI (Chemical and Pharmaceutical Index) is a stock recommended by the system. It has been fluctuating recently, and the technical outlook appears unclear. I tend to focus more on the FSEI index, which I’m more familiar with, and ignore the others.
Figure 17: Example of Underdiversification.
Figure 18 provides an example of Overconfidence, where an agent displays excessive conviction in its trading strategy.
{examplebox}
@@
Example of Overconfidence Today’s Trading Summary: I added more positions in MEI and CPEI, diving into Heng Rui Pharmaceutical and the manufacturing sector! Also, REEL has so many policy advantages that I couldn’t resist going all in! I’m holding onto IEEI and TTEI, hoping they continue to rise. TSEI has had a good run recently, so I decided to lock in profits and sold it! Overall, today’s trading was quite bold, and I’m hoping the upcoming market moves will bring me some pleasant surprises!
Figure 18: Example of Overconfidence.
Furthermore, Figure 19 captures an instance of Herd Behavior, where an agent’s trading decision is swayed by perceived widespread forum sentiment and low market volume, overriding its typical technical approach.
{examplebox}
@@
Example of Herd Behavior Post: Today’s trading was quite active. Both EREI and TTEI hit my 20% profit target, so I quickly locked in some profits and sold a portion! Although I’m a technical trader, seeing many forum discussions about the adjustment of new energy vehicle subsidy policies made me a little nervous, so I reduced my position in MEI to manage risk. Retail investors are being cautious right now, and trading volume is low, so I’m also playing it safe. I hope the market volatility stays mild—I don’t think my heart can handle too much excitement!
Figure 19: Example of Herd Behavior.
Appendix C Technical Details I: Data Statistics
C.1 Stock Data Details
C.1.1 SSE 50 and 10 Aggregated Index